OpenStudy (anonymous):

solve for x in the proportion (3x)/(3x-36)= -1/x

OpenStudy (precal):

cross multiply

OpenStudy (anonymous):

ik but i can't find the answer plz help

OpenStudy (anonymous):

@precal help

OpenStudy (precal):

@nicole1217 sorry I have been busy, let me help you

OpenStudy (precal):

\[\frac{3x}{3x-36}=\frac{-1}{x}\] cross multiply \[3x^2=-1(3x-36)\] \[3x^2=-3x+36\] \[3x^2+3x-36=0\]

OpenStudy (precal):

\[3(x^2+x-12)=0\] 3(x+4)(x-3)=0 x=-4 x=3